# LLM4DyG: Can Large Language Models Solve Spatial-Temporal Problems on Dynamic Graphs?

### v2

@@ Abstract.
In an era marked by the increasing adoption of Large Language Models (LLMs) for various tasks, there is a growing focus on exploring LLMs’ capabilities in handling web data, particularly graph data. Dynamic graphs, which capture temporal network evolution patterns, are ubiquitous in real-world web data. Evaluating LLMs’ competence in understanding spatial-temporal information on dynamic graphs is essential for their adoption in web applications, which remains unexplored in the literature. In this paper, we bridge the gap via proposing to evaluate LLMs’ spatial-temporal understanding abilities on dynamic graphs, to the best of our knowledge, for the first time. Specifically, we propose the LLM4DyG benchmark, which includes nine specially designed tasks considering the capability evaluation of LLMs from both temporal and spatial dimensions. Then, we conduct extensive experiments to analyze the impacts of different data generators, data statistics, prompting techniques, and LLMs on the model performance. Finally, we propose Disentangled Spatial-Temporal Thoughts (DST2) for LLMs on dynamic graphs to enhance LLMs’ spatial-temporal understanding abilities. Our main observations are: 1) LLMs have preliminary spatial-temporal understanding abilities on dynamic graphs, 2) Dynamic graph tasks show increasing difficulties for LLMs as the graph size and density increase, while not sensitive to the time span and data generation mechanism, 3) the proposed DST2 prompting method can help to improve LLMs’ spatial-temporal understanding abilities on dynamic graphs for most tasks. The data and codes will be open-sourced at publication time.
@@ 1. Introduction
In an era marked by the increasing adoption of Large Language Models (LLMs) for various tasks beyond natural language processing, such as image recognition (Alayrac et al., 2022), healthcare diagnostics (Thirunavukarasu et al., 2023), and autonomous agents (Wang et al., 2023b), there has been a growing body of research dedicated to exploring LLMs’ abilities to tackle the vast troves of web data. One area of particular interest is the handling of graph data, which ubiquitously exists on the Internet. The World Wide Web itself can be seen as a colossal interconnected graph of webpages, hyperlinks, and content. For example, social media platforms like Facebook, Twitter, and Instagram generate dynamic social graphs reflecting user interactions and connections.
To leverage the in-context learning and commonsense knowledge of LLMs, several pioneer works have been dedicated to adopting LLMs on static graphs. For instance,  Wang et al. (2023a) and  Guo et al. (2023) propose benchmarks to evaluate LLMs’ proficiency in comprehending and reasoning about graph structures, with tasks like graph connectivity, topological sort, etc, demonstrating the LLMs’ abilities of in-context learning and reasoning to solve static graph problems.  Ye et al. (2023) and  Chen et al. (2023b) propose to fine-tune the LLMs to solve graph tasks in natural language, showing the strong potential of LLMs to leverage text information, generate human-readable explanations, and integrate commonsense knowledge to enhance the reasoning over structures.
@@
To this end, in this paper we propose to explore the following research question, Can LLMs understand and handle the spatial-temporal information on dynamic graphs in natural language?
However, this problem remains unexplored in literature, and is non-trivial with the following challenges: 1) How to design dynamic graph tasks to assess the capabilities of LLMs to understand temporal and structural information both separately and simultaneously; 2) How to investigate the impacts of spatial and temporal dimensions, where they have complex and mixed interactions on dynamic graphs; 3) How to design the prompts for dynamic graphs and tasks, where spatial-temporal information should be taken into consideration in natural language.
@@ 2.1. LLMs for tasks with graph data
Recently, there has been a surge of works about LLMs for solving tasks with graph data (Zhang et al., 2023). He et al. (2023) proposed an approach that LLMs not only execute zero-shot predictions but also generate coherent explanations for their decisions. These explanations are subsequently leveraged to enhance the features of graph nodes for node classification in text-attributed graphs.  Chen et al. (2023b) proposes to explore LLMs-as-Enhancers and LLMs-as-Predictors for solving graph-related tasks, where the former augment the GNN with LLMs, and the latter directly adopts LLMs to make predictions. Wang et al. (2023a) introduced NLGraph, a benchmarking framework tailored for evaluating the performance of LLMs on traditional graph-related tasks. Simultaneously,  Guo et al. (2023) conducted a comprehensive empirical study focused on utilizing LLMs to tackle structural and semantic understanding tasks within graph-based contexts. Recent contributions in this line of research include InstructGLM (Ye et al., 2023), a method for fine-tuning LLMs inspired by LLaMA (Touvron et al., 2023), designed specifically for node classification tasks.  Zhang (2023) and  Jiang et al. (2023) have initiated the exploration of this frontier by interfacing LLMs with external tools and enhancing their reasoning capabilities over structured data sources such as knowledge graphs (KGs) and tables. However, these works mainly focus on static graphs, ignoring the temporal nature of graphs in real-world web applications. In this paper, we propose to explore LLMs on dynamic graphs, which remains unexplored in the literature.
@@ 2.2. LLMs for other related tasks
LLMs have been recently applied to other related tasks, including time-series forecasting, recommendation, etc.  Yu et al. (2023) presents a novel study on harnessing LLMs’ outstanding knowledge and reasoning abilities for explainable financial time series forecasting. Chang et al. (2023) leverages pre-trained LLMs to enhance time-series forecasting and has shown exceptional capabilities as both a robust representation learner and an effective few-shot learner.  Sun et al. (2023) summarizes two strategies for completing time-series (TS) tasks using LLM: LLM-for-TS that designs and trains a fundamental large model for TS data and TS-for-LLM that enables the pre-trained LLM to handle TS data.  Lyu et al. (2023) investigates various prompting strategies for enhancing personalized recommendation performance with large language models through input augmentation. However, these works do not consider the role of structures, and in this paper, we mainly focus on exploring the spatial-temporal understanding abilities of LLMs on dynamic graphs.
@@ 2.3. Dynamic Graph Learning
Dynamic graphs are pervasive in a multitude of real-world applications, spanning areas such as event forecasting, recommendation systems, and many more (Cai et al., 2021; Deng et al., 2020; You et al., 2019; Wang et al., 2021b; Li et al., 2019; Wu et al., 2020a). This prevalence has prompted significant research interest in the development and refinement of dynamic graph neural networks (Skarding et al., 2021; Zhu et al., 2022; Chen et al., 2023a). These networks are designed to model intricate graph dynamics, which incorporate evolving structures and features over time. A variety of approaches have been proposed to address the challenges posed by dynamic graphs. Some research efforts have focused on employing Graph Neural Networks (GNNs) (Zhou et al., 2020; Wu et al., 2020b; Wang et al., 2019, 2017; Xu et al., 2018) to aggregate neighborhood information for each individual snapshot of the graph. Subsequently, these methods use a sequence module to capture and model the temporal information (Yang et al., 2021; Sun et al., 2021; Hajiramezanali et al., 2019; Seo et al., 2018; Sankar et al., 2020). In contrast, other studies have proposed the use of time-encoding techniques. These methods encode the temporal links into specific time-aware embeddings, and then utilize a GNN or memory module (Wang et al., 2021a; Cong et al., 2021; Xu et al., 2020; Rossi et al., 2020) to process and handle the structural information embedded in the graph. However, these methods require the model to be trained every time they encounter a new dynamic graph task, limiting their widespread usage in real-world scenarios. In this paper, we explore the potential of LLMs on solving dynamic graph tasks with in-context learning skills and evaluate their spatial-temporal understanding abilities on dynamic graphs.
@@ 3.1. Dynamic Graph Data Generators
We first adopt a random dynamic graph data generator to control the statistics of the dynamic graph. In default, we adopt an Erdős-Rényi (ER) model to generate an undirected graph, and randomly assign a time-stamp for each edge. Denote a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with the node set $\mathcal{V}=\{v_{1},v_{2},\dots v_{N}\}$ and edge set =$\mathcal{E}=\{e_{1},e_{2},\dots,e_{M}\}$. We first generate the graph with the ER model $\mathcal{G}=ER(N,p)$ where $N$ is the number of nodes in the graph, and $p$ is the probability of edge occurrence between each node pair. In this way, $N$ controls the graph size, and $p$ controls the graph density. After obtaining the graph $\mathcal{G}$, we assign each edge with a random timestamp $t\sim U(\{0,1,\dots,T-1\})$, where $T$ controls the time span. For a generated dynamic graph, each edge $e=(v_{i},v_{j},t)$ denotes that node $v_{i}$ and node $v_{j}$ are linked at time $t$. We also include other dynamic graph generators, stochastic block (SB) model, and forest fire (FF) model.
@@ 4. Experiments
Prompt
Example
DyG Instruction
In an undirected dynamic graph, (u, v, t) means that node u and node v are linked with an undirected edge at time t.
Task Instruction
Your task is to answer when two nodes are first connected in the dynamic graph. Two nodes are connected if there exists a path between them.
Answer Instruction
Give the answer as an integer number at the last of your response after ’Answer:’
Exemplar
Here is an example: Question: Given an undirected dynamic graph with the edges [(0, 1, 0), (1, 2, 1), (0, 2, 2)]. When are node 0 and node 2 first connected? Answer:1
Question
Question: Given an undirected dynamic graph with the edges [(0, 9, 0), (1, 9, 0), (2, 5, 0), (1, 2, 1), (2, 6, 1), (3, 7, 1), (4, 5, 2), (4, 7, 2), (7, 8, 2), (0, 1, 3), (1, 6, 3), (5, 6, 3), (0, 4, 4), (3, 4, 4), (3, 6, 4), (4, 6, 4), (4, 9, 4), (6, 7, 4)]. When are node 2 and node 1 first connected?
Answer
Answer:1
@@ 4.4. Results with different LLMs
Generation Model ER Model SB Model FF Model
zero-shot 2.3$\pm$0.5 7.7$\pm$1.7 5.3$\pm$2.5
one-shot 33.7$\pm$2.1 46.0$\pm$2.9 48.0$\pm$7.1
zero-shot COT 1.0$\pm$0.8 5.7$\pm$3.1 2.0$\pm$1.6
one-shot COT 10.3$\pm$0.5 15.3$\pm$0.9 13.0$\pm$2.9
@@ Appendix A Additional Experiments
We add experiments on three larger-scale real-world datasets, which are described as follows
@@
We compare the performance (ACC) of static baseline methods NLGraph (Wang et al., 2023a), GPT4Graph (Guo et al., 2023) in the spatial task ’neighbors at time’. As shown in Table 8, our method has a significant performance over the baselines on larger-scale real-world dynamic graphs, which can be credited to our consideration of temporal information on dynamic graphs. Though we focus on evaluating LLMs’ spatial-temporal understanding abilities on dynamic graphs, the proposed method out-performs the recent baselines for static graphs in real-world datasets.
@@
we also compare the static baselines (Wang et al., 2023a; Guo et al., 2023) on spatial tasks. As shown in Table 9, our method out-performs the recent static baselines, showing the effectiveness of our method for tackling problems on dynamic graphs.
Table 9. Comparisons (ACC) with static methods on spatial tasks.
@@
Table 10. Comparisons with static methods on link prediction tasks.
@@
we add experiments for link prediction tasks to further compare the baselines (Wang et al., 2023a; Guo et al., 2023) for predictive tasks. As shown in Table 10, our method out-performs static baselines NLGraph and GPT4Graph with a large margin, demonstrating the effectiveness of handling spatial-temporal information on dynamic graphs.

### v3

@@ Abstract.
In an era marked by the increasing adoption of Large Language Models (LLMs) for various tasks, there is a growing focus on exploring LLMs’ capabilities in handling web data, particularly graph data. Dynamic graphs, which capture temporal network evolution patterns, are ubiquitous in real-world web data. Evaluating LLMs’ competence in understanding spatial-temporal information on dynamic graphs is essential for their adoption in web applications, which remains unexplored in the literature. In this paper, we bridge the gap via proposing to evaluate LLMs’ spatial-temporal understanding abilities on dynamic graphs, to the best of our knowledge, for the first time. Specifically, we propose the LLM4DyG benchmark, which includes nine specially designed tasks considering the capability evaluation of LLMs from both temporal and spatial dimensions. Then, we conduct extensive experiments to analyze the impacts of different data generators, data statistics, prompting techniques, and LLMs on the model performance. Finally, we propose Disentangled Spatial-Temporal Thoughts (DST2) for LLMs on dynamic graphs to enhance LLMs’ spatial-temporal understanding abilities. Our main observations are: 1) LLMs have preliminary spatial-temporal understanding abilities on dynamic graphs, 2) Dynamic graph tasks show increasing difficulties for LLMs as the graph size and density increase, while not sensitive to the time span and data generation mechanism, 3) the proposed DST2 prompting method can help to improve LLMs’ spatial-temporal understanding abilities on dynamic graphs for most tasks. The data and codes are publicly available at Github.
Dynamic Graph; Large Language Model; Spatial-Temporal; Benchmark; Evaluation; Disentanglement
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671709isbn: 979-8-4007-0490-1/24/08ccs: Information systems Data miningccs: Computing methodologies Natural language processingccs: Computing methodologies Knowledge representation and reasoning
@@ 1. Introduction
In an era marked by the increasing adoption of Large Language Models (LLMs) for various tasks beyond natural language processing, such as image recognition (Alayrac et al., 2022), healthcare diagnostics (Thirunavukarasu et al., 2023), and autonomous agents (Wang et al., 2023c), there has been a growing body of research dedicated to exploring LLMs’ abilities to tackle the vast troves of web data. One area of particular interest is the handling of graph data, which ubiquitously exists on the Internet. The World Wide Web itself can be seen as a colossal interconnected graph of webpages, hyperlinks, and content. For example, social media platforms like Facebook, Twitter, and Instagram generate dynamic social graphs reflecting user interactions and connections.
To leverage the in-context learning and commonsense knowledge of LLMs, several pioneer works have been dedicated to adopting LLMs on static graphs. For instance,  Wang et al. (2023b) and  Guo et al. (2023) propose benchmarks to evaluate LLMs’ proficiency in comprehending and reasoning about graph structures, with tasks like graph connectivity, topological sort, etc, demonstrating the LLMs’ abilities of in-context learning and reasoning to solve static graph problems.  Ye et al. (2023) and  Chen et al. (2023b) propose to fine-tune the LLMs to solve graph tasks in natural language, showing the strong potential of LLMs to leverage text information, generate human-readable explanations, and integrate commonsense knowledge to enhance the reasoning over structures.
@@
To this end, in this paper, we propose to explore the following research question:
Can large language models understand and handle the spatial-temporal information on dynamic graphs in natural language?
However, this problem remains unexplored in literature, and is non-trivial with the following challenges:
  • How to design dynamic graph tasks to assess the capabilities of LLMs to understand temporal and structural information both separately and simultaneously.
  • How to investigate the impacts of spatial and temporal dimensions, where they have complex and mixed interactions on dynamic graphs.
  • How to design the prompts for dynamic graphs and tasks, where spatial-temporal information should be taken into consideration in natural language.
@@ 2.1. LLMs for tasks with graph data
Recently, there has been a surge of works about LLMs for solving tasks with graph data (Zhang et al., 2023c, b; Li et al., 2023; Jin et al., 2023). He et al. (2023) proposed an approach that LLMs not only execute zero-shot predictions but also generate coherent explanations for their decisions. These explanations are subsequently leveraged to enhance the features of graph nodes for node classification in text-attributed graphs.  Chen et al. (2023b) proposes to explore LLMs-as-Enhancers and LLMs-as-Predictors for solving graph-related tasks, where the former augment the GNN with LLMs, and the latter directly adopts LLMs to make predictions. Wang et al. (2023b) introduced NLGraph, a benchmarking framework tailored for evaluating the performance of LLMs on traditional graph-related tasks. Simultaneously,  Guo et al. (2023) conducted a comprehensive empirical study focused on utilizing LLMs to tackle structural and semantic understanding tasks within graph-based contexts. Recent contributions in this line of research include InstructGLM (Ye et al., 2023), a method for fine-tuning LLMs inspired by LLaMA (Touvron et al., 2023), designed specifically for node classification tasks.  Zhang (2023) and  Jiang et al. (2023b) have initiated the exploration of this frontier by interfacing LLMs with external tools and enhancing their reasoning capabilities over structured data sources such as knowledge graphs (KGs) and tables.  Yao et al. (2024) explores leveraging LLMs for graph generation, which have potentials for various real-world tasks like drug discoveries.  Li et al. (2024b) utilizes LLMs to recognize and provide the possible latent causal structures given the rich domain knowledge in the textual corpus. However, these works mainly focus on static graphs, ignoring the temporal nature of graphs in real-world web applications. In this paper, we propose to explore LLMs’ spatial-temporal understanding on dynamic graphs, which remains unexplored in the literature.
@@ 2.2. LLMs for other related tasks
LLMs have been recently applied to other related tasks, including time-series forecasting, recommendation, etc.  Yu et al. (2023) presents a novel study on harnessing LLMs’ outstanding knowledge and reasoning abilities for explainable financial time series forecasting. Chang et al. (2023) leverages pre-trained LLMs to enhance time-series forecasting and has shown exceptional capabilities as both a robust representation learner and an effective few-shot learner.  Sun et al. (2023) summarizes two strategies for completing time-series (TS) tasks using LLM: LLM-for-TS that designs and trains a fundamental large model for TS data and TS-for-LLM that enables the pre-trained LLM to handle TS data.  Feng et al. (2023) and  (Yipeng Zhang, 2024) evaluate the temporal or sequential understanding abilities of LLMs in visual tasks.  Lyu et al. (2023) investigates various prompting strategies for enhancing personalized recommendation performance with large language models through input augmentation. However, these works do not consider the role of structures, and in this paper, we mainly focus on exploring the spatial-temporal understanding abilities of LLMs on dynamic graphs.
@@ 2.3. Dynamic Graph Learning
Dynamic graphs are pervasive in a multitude of real-world applications, spanning areas such as event forecasting, recommendation systems, etc (Cai et al., 2021; Deng et al., 2020; You et al., 2019; Wang et al., 2021b; Li et al., 2019; Wu et al., 2020). This prevalence has prompted significant research interest in the development and refinement of dynamic graph neural networks (Skarding et al., 2021; Zhu et al., 2022; Chen et al., 2023a; Zhang et al., 2023h). These networks are designed to model intricate graph dynamics, which incorporate evolving structures and features over time. A variety of approaches have been proposed to address the challenges posed by dynamic graphs. Some research efforts have focused on employing Graph Neural Networks (GNNs) to aggregate neighborhood information for each individual snapshot of the graph. Subsequently, these methods use a sequence module to capture and model the temporal information (Yang et al., 2021; Sun et al., 2021; Hajiramezanali et al., 2019; Seo et al., 2018; Sankar et al., 2020). In contrast, other studies have proposed the use of time-encoding techniques. These methods encode the temporal links into specific time-aware embeddings, and then utilize a GNN or memory module (Wang et al., 2021a; Cong et al., 2021; Xu et al., 2020; Rossi et al., 2020) to process and handle the structural information embedded in the graph. Recently, there are some works focusing on studying dynamic graphs under distribution shifts (Zhang et al., 2023a, d, e, 2022a). However, these methods require the model to be trained every time they encounter a new dynamic graph task, limiting their widespread usage in real-world scenarios. In this paper, we explore the potential of LLMs on solving dynamic graph tasks with in-context learning skills and evaluate their spatial-temporal understanding abilities on dynamic graphs.
@@ 3.1. Dynamic Graph Data Generators
We first adopt a random dynamic graph data generator to control the statistics of the dynamic graph. In default, we adopt an Erdős-Rényi (ER) model to generate an undirected graph, and randomly assign a time-stamp for each edge. Denote a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with the node set $\mathcal{V}=\{v_{1},v_{2},\dots v_{N}\}$ and edge set =$\mathcal{E}=\{e_{1},e_{2},\dots,e_{M}\}$. We first generate the graph with the ER model $\mathcal{G}=ER(N,p)$ where $N$ is the number of nodes in the graph, and $p$ is the probability of edge occurrence between each node pair. In this way, $N$ controls the graph size, and $p$ controls the graph density. After obtaining the graph $\mathcal{G}$, we assign each edge with a random timestamp $t\sim U(\{0,1,\dots,T-1\})$, where $T$ controls the time span. For a generated dynamic graph, each edge $e=(v_{i},v_{j},t)$ denotes that node $v_{i}$ and node $v_{j}$ are linked at time $t$. We also include other dynamic graph generators, stochastic block (SB) model, and forest fire (FF) model. For real-world datasets, we adopt a random sampler to sample ego-graphs from the real graphs for evaluation.
@@ 4. Experiments
Prompt Example
DyG Instruction In an undirected dynamic graph, (u, v, t) means that node u and node v are linked with an undirected edge at time t.
Task Instruction Your task is to answer when two nodes are first connected in the dynamic graph. Two nodes are connected if there exists a path between them.
Answer Instruction Give the answer as an integer number at the last of your response after ’Answer:’
Exemplar Here is an example: Question: Given an undirected dynamic graph with the edges [(0, 1, 0), (1, 2, 1), (0, 2, 2)]. When are node 0 and node 2 first connected? Answer:1
Question Question: Given an undirected dynamic graph with the edges [(0, 9, 0), (1, 9, 0), (2, 5, 0), (1, 2, 1), (2, 6, 1), (3, 7, 1), (4, 5, 2), (4, 7, 2), (7, 8, 2), (0, 1, 3), (1, 6, 3), (5, 6, 3), (0, 4, 4), (3, 4, 4), (3, 6, 4), (4, 6, 4), (4, 9, 4), (6, 7, 4)]. When are node 2 and node 1 first connected?
Answer Answer:1
@@ 4.4. Results with different LLMs
Generation Model ER Model SB Model FF Model
zero-shot 2.3$\pm$0.5 7.7$\pm$1.7 5.3$\pm$2.5
one-shot 33.7$\pm$2.1 46.0$\pm$2.9 48.0$\pm$7.1
zero-shot COT 1.0$\pm$0.8 5.7$\pm$3.1 2.0$\pm$1.6
one-shot COT 10.3$\pm$0.5 15.3$\pm$0.9 13.0$\pm$2.9
@@ 5. Conclusion
Acknowledgements.
This work is supported by the National Key Research and Development Program of China No. 2023YFF1205001, National Natural Science Foundation of China (No. 62222209, 62250008, 62102222, 62206149), Beijing Key Lab of Networked Multimedia and Beijing National Research Center for Information Science and Technology under Grant No. BNR2023RC01003, BNR2023TD03006.
@@ Appendix A Additional Experiments
We add experiments on three larger-scale real-world datasets. The details are as follows:
@@
We compare the performance (ACC) of static baseline methods NLGraph (Wang et al., 2023b), GPT4Graph (Guo et al., 2023) in the spatial task ’neighbors at time’. As shown in Table 8, our method has a significant performance over the baselines on larger-scale real-world dynamic graphs, which can be credited to our consideration of temporal information on dynamic graphs. Though we focus on evaluating LLMs’ spatial-temporal understanding abilities on dynamic graphs, the proposed method out-performs the recent baselines for static graphs in real-world datasets.
Table 7. Task instructions in DyG prompts.
Task Task Instruction
when link Your task is to answer when two nodes are linked in the dynamic graph.
when   connect Your task is to answer when two nodes are first connected in the dynamic graph. Two nodes are connected if there exists a path between them.
when   tclosure Your task is to answer when the three nodes in the dynamic graph first close the triad. Two nodes with a common neighbor is said to have a triadic closure, if they are linked since some time so that the three nodes have linked with each other to form a triad.
neighbor at time Your task is to answer what nodes are linked with a given node at a given time in the dynamic graph.
neighbor in periods Your task is to answer what nodes are linked with one node only after some time in the dynamic graph
check tclosure Your task is to answer whether three nodes in the dynamic graph formed a closed triad. A closed triad is composed of three nodes which have linked with each other some time.
check tpath Your task is to answer whether a path is chronological in the dynamic graph. The time of the edges in a chronological path from source node to target node must not decrease, e.g., [2, 3, 5] is a chronological path in the dynamic graph [(2, 3, 0), (3, 5, 1)]
find tpath Your task is to find a chronological path in the dynamic graph. The time of the edges in a chronological path from source node to target node must not decrease, e.g., [2, 3, 5] is a chronological path in the dynamic graph [(2, 3, 0), (3, 5, 1)]
sort edge Your task is to sort the edges in the dynamic graph by time from earlest to latest.
@@
We compare the static baselines (Wang et al., 2023b; Guo et al., 2023) on spatial tasks. As shown in Table 9, our method out-performs the recent static baselines, showing the effectiveness of our method for tackling problems on dynamic graphs.
Table 9. Comparisons (ACC) on several spatial tasks.
@@
Table 10. Comparisons on link prediction tasks.
@@
we add experiments for link prediction tasks to further compare the baselines (Wang et al., 2023b; Guo et al., 2023) for predictive tasks. As shown in Table 10, our method out-performs static baselines NLGraph and GPT4Graph with a large margin, demonstrating the effectiveness of handling spatial-temporal information on dynamic graphs.
Table 11. Comparisons (ACC) with different time formatting types on several tasks.
Time types when tclosure check tclosure
Original 0.59 0.56
UNIX 0.35 0.53
Date 0.47 0.57
Effects of time formatting types.
We add experiments with different types of timestamps, including original, UNIX timestamp, Date, using Codellama2-13B and two tasks. The results are shown in the table 11. The original timestamp is an integer ranging from 1 to $T$, where T is the time span. The UNIX timestamp ranges from 2010-01-01 00:00:00 to 2020-01-01 00:00:00. The date is formatted as ‘%Year%Month%Day’, e.g., ‘20200101’. The results show that using UNIX timestamp or date for time formatting reduces the performance, which may be due to the increased complexity for LLMs to infer the ordering between time.
Table 12. Comparisons (ACC) with different node formatting types on several tasks.
Node name types when tclosure check tclosure
Original 0.59 0.56
Random indexes 0.62 0.55
People names 0.68 0.62
Effects of node formatting types.
We add experiments with different types of node names, including original, random indexes, people names, using Codellama2-13B and two tasks. The results are shown in the table 12. The original node formatting uses integers ranging from 0 to $N$ to represent the nodes, where $N$ refers to the number of nodes, while the random indexes adopt integers ranging from 0 to 1e8. For the formatting of people names, several names are adopted to represent nodes, e.g., ’Aiden’, ’Priya’, ’Dmitri’. One interesting phenomenon is that using names to represent nodes may improve the performance, which may be due to that LLMs may be more familiar with names than integers. The results are consistent with previous literature(Fatemi et al., 2024).